OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $2,000, with an annual interest rate of 18 percent and a term of 3 years. Round to the nearest cent as needed

OpenStudy (chaise):

I will assume that this is simple interest. A=PiT A=2,000 * 0.18 * 3 = 1080. $1080 will be paid in interest, plus the original loan of $2000. 1080+2000=3080. 3080 must be paid in 3 years. There are 12 months in a year. 36 months in 3 years 3080/3=85 and 56 cents. each month $85 and 56 cents will need to be repaid.

OpenStudy (kropot72):

The amount to be paid off assuming the interest compounds monthly is: \[2000(1+(.18/12))^{36}=2000\times1.70914=$3,418.28\] The monthly payment is: \[3418.28\div36=$94.95\]

OpenStudy (anonymous):

kropot72 -- that assumes no payments until the end. With equal monthly payments, it should be amortized. I think the answer by chaise is what was wanted.